Consider the following set of data.

(19, 11), (26, 53), (65, 28), (77, 20), (108, 56), (118, 11)

(a) Calculate the covariance of the set of data. (Give your answer correct to two decimal places.)
A = -5.63

(b) Calculate the standard deviation of the six x-values and the standard deviation of the six y-values. (Give your answers correct to three decimal places.)

sx =  
sy =

(c) Calculate r, the coefficient of linear correlation, for the data in part (a). (Give your answer correct to two decimal places.)
Answer #1

ANSWER:-

X Y X-ar{X} Y-ar{Y} (X - X)(Y- Y)
19 11 ( 19-68.83)--49.83 (11-29.83)--18.33 (-49.83 *-18.83) 938.2989
26 53 (26-68.83)=-42.83 (53-29.83)=23.17 (-42.83 23.17)-992.3711
65 28 (65-68.83)=-3.83 (28-29.83)=-1.83 (-3..83*-1.83)=7.0089
77 20 (77-68.83)=8.17 (20-29.83)=-9.83 (8.17*-9.83)=-80.3111
108 56 (108-68.83)=39.17 (56-29.83)=26.17 (39.17*26.17)=1025.0789
118 11 (118-68.83)=49.17 (11-29.83)--18.33 (49.17*-18.83)=-925.8711

ar{X}=rac{19+26+65+77+108+118}{6}

ar{X}=rac{413}{6}=68.8333

ar{X}=68.83

ar{Y}=rac{11+53+28+20+56+11}{6}

ar{Y}=rac{179}{6}

ar{Y}=29.83
calculation of Cov(X,Y):-

We know that

Cov(X,Y)=rac{sum (X-ar{X})(Y-ar{Y})}{n-1}

Cov(X,Y)=rac{938.2989-992.3711+7.0089-80.3111+1025.0789-925.8711}{5}

Cov(X,Y)=rac{-28.1666}{5}

Cor(Х. У)--5.633

Calculation of Sx:-

Sx=sqrt{rac{sum (X-ar{X})^2}{n}}

Sx=sqrt{rac{(-49.83)^2+(-42.83)^2+(-3.83)^2+(8.17)^2+(39.17)^2+(49.17)^2)}{6}}

Sx=sqrt{rac{2483.0289+1834.4089+14.6689+66.7489+1534.2889+2417.6889}{6}}

8350.8334 s-V

Sx=sqrt{1391.805567}

Sx=37.3069

Sx=37.307

Calculation of Sy:-

Sy=sqrt{rac{sum (Y-ar{Y})^2}{n}}

Sy=sqrt{rac{(18.83)^2+(23.17)^2+(-1.83)^2+(-9.83)^2+(26.17)^2+(-18.83)^2)}{6}}

354.5689 + 536.8489 + 3.3489 96.6289 +684.8689 + 354.5689 6 Sy

2030.8334 6

Sy=sqrt{338.4722}

Sy 18.3976

Sy=18.398

calculationo fr:

We know that

r=rac{Cov(X,Y)}{SxSy}

Where,

Cor(Х. У)--5.633

Sx=37.307

Sy=18.398

8.2074 103

r=-0.00821